OpenStudy (anonymous):

Help plz!! Find the standard form of the equation of the parabola with a focus at (3, 0) and a directrix at x = -3.

OpenStudy (anonymous):

y = one divided by twelvex2 -12y = x2 x = one divided by twelvey2 y2 = 6x

OpenStudy (anonymous):

Those are the possible answers and I think its the third one I just need clarification

OpenStudy (tkhunny):

"I think its the" - Nope. Never good enough. Prove it! Definition of Parabola. Set of points equidistant from a given point (focus) and a given line (directrix). Given Point: (3,0) Given Line: x = -3 What is the distance of ANY point, (x,y) from (3,0)? What is the distance of ANY point, (x,y) from x = -3?

OpenStudy (anonymous):

3 and -3 I'm assuming the (x,y) is at the origin? (totally clueless)

OpenStudy (tkhunny):

I said ANY point, not the Origin. ANY point, (x,y) is this far from (3,0), \(\sqrt{(x-3)^{2}+(y-0)^{2}}\). The distance formula. Are we ringing any bells?

OpenStudy (anonymous):

Oh! I know the formula but to use it for this wasn't in the lesson.

OpenStudy (anonymous):

ok so from that equation I cant really get an answer so do I plug in values for the x and y?

OpenStudy (tkhunny):

If you know it, it's for this lesson. Remember EVERYTHING. Okay, what is the distance of ANY point, (x,y) from x = -3? Try a few points and see.

OpenStudy (anonymous):

Well I plugged in (1,1) into the equation and got Sqrt 5 Did it again from (2,7) and got 5Sqrt2

OpenStudy (tkhunny):

No, we're no longer worried about the distance from (3,0). We have that, already. It's that lovely square root from the distance formula. NOW, we need to know the distance from x = -3.

OpenStudy (anonymous):

So what formula asre yu saying I should be using?

OpenStudy (anonymous):

Are you saying to use slope or the same distance formula?

OpenStudy (anonymous):

If we are going from the focus then the distance would be 6

OpenStudy (tkhunny):

Geometry. Draw a perpendicular line from x = -3 through the desired point. Measure the distance. The point (0,4) is a distance of 3 from x = -3. The point (1,3) is a distance of 4 from x = -3. The point (-2,2) is a distance of 1 from x = -3. You have to see it!

OpenStudy (anonymous):

I see that. Sorry didnt understand you meant from random points again.

OpenStudy (tkhunny):

They are NOT random points. We are talking about EVERY CONCEIVABLE point.

OpenStudy (tkhunny):

Given a point, I don't care which one, how far is it from the focus? Given that same point, whichever one we picked in the previous quiz, how far is it from x = -3?

OpenStudy (anonymous):

Ok so from (0,4) the distance would be 3 and from it would be 6 from the focus

OpenStudy (anonymous):

*and it would be 6 from the focus

OpenStudy (tkhunny):

We don't care about (0,4). That is just one sample point from infinitely many points. We have to get them ALL. All points are this far from (3,0): \(\sqrt{(x-3)^{2}+(y)^{2}}\) All points are this far from x = -3: \(|x-(-3)| = |x+3|\) You must see these two things. You can pick ANY point and calculate those two distances. However, we don't have a parabola until we do this: \(\sqrt{(x-3)^{2}+y^{2}} = |x+3|\) Now THAT is a parabola. Your task is to put it in standard form.

OpenStudy (anonymous):

oh ok!! Yay! So for problems like these in the future do I just follow the set up that we just went through?

OpenStudy (tkhunny):

I hope not. We should go directly to the solution, rather than walking around the block a few times. I hope not. This is the focus-directrix DEFINITION of a parabola. It is useful mostly for proving existence. Definitions are often cumbersome therefore, not particularly helpful for more common use. I hope not. You should learn a couple other methods along the way. Keep your eyes and ears open and understand what you are doing. Don't try to memorize some process.

OpenStudy (anonymous):

Okay, thank you so much! I got x=1/12y^2 btw :)

OpenStudy (tkhunny):

Is that right?

OpenStudy (anonymous):

Yep! :)

OpenStudy (tkhunny):

Perfect.